OpenStudy (anonymous):

Imagine that you took a road trip. Based on the information in the table, what was the average speed of your car? Time Mile marker 3:00 pm 11 8:00 pm 205

OpenStudy (anonymous):

Speed is distance covered over time (even though this is velocity, it still applies, in this case): \[ v=\frac{\Delta x}{\Delta t}=\frac{205-11}{8-3}=\frac{194}{5} \] And we are done.

OpenStudy (anonymous):

wow cool

OpenStudy (anonymous):

and units?

OpenStudy (anonymous):

Oh, well, we're dealing with how many miles recurred in a set of hours, so, mph.